# Tidal effects based on GUP-induced effective metric

### v1

@@ V Geodesic deviation equations of Schwarzschild black hole with effective metric
On the other hand, in Fig. 7, we have drawn the angular components of the geodesic deviation in the Schwarzschild black hole modified by the $\alpha$ dependent effective metric. Fig. 7(a) corresponds to $\alpha=0$ with zero initial velocity, $\frac{d\eta^{\hat{1}}(b)}{d\tau}=0$, the original Schwarzschild case. When the initial velocity of the original Schwarzschild black hole is non-zero as $\frac{d\eta^{\hat{1}}(b)}{d\tau}\neq 0$, the angular separation is drawn by the dashed curve in Fig. 7(b). Now, the angular components of the geodesic deviation in the Schwarzschild black hole modified by an effective metric by varying the GUP parameter $\alpha$ are drawn by the solid curves in Fig. 7(b). Note that as the GUP parameters $\alpha$ increase, the angular separation appears to be larger in all ranges except near the singularity, where the angular separation is inverted to be small. Moreover, in Fig. 8, we have also drawn the angular components of the geodesic deviation by varying the mass. The original Schwarzschild case is drawn in Fig. 8(a). The GUP-modified Schwarzschild black hole case is drawn in Fig. 8(b), which shows that it resembles the original Schwarzschild case but differs near the singularities.

### v2

@@ V Geodesic deviation equations of Schwarzschild black hole with effective metric
On the other hand, in Fig. 7, we have drawn the angular components of the geodesic deviation in the Schwarzschild black hole modified by the $\alpha$ dependent effective metric. Fig. 7(a) corresponds to $\alpha=0$ with zero initial velocity, $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}=0$, the original Schwarzschild case. When the initial velocity of the original Schwarzschild black hole is non-zero as $\frac{\rm{d}\eta^{\hat{1}}(b)}{\rm{d}\tau}\neq 0$, the angular separation is drawn by the dashed curve in Fig. 7(b). Now, the angular components of the geodesic deviation in the Schwarzschild black hole modified by an effective metric by varying the GUP parameter $\alpha$ are drawn by the solid curves in Fig. 7(b). Note that as the GUP parameters $\alpha$ increase, the angular separation appears to be larger in all ranges except near the singularity, where the angular separation is inverted to be small. Moreover, in Fig. 8, we have also drawn the angular components of the geodesic deviation by varying the mass. The original Schwarzschild case is drawn in Fig. 8(a). The GUP-modified Schwarzschild black hole case is drawn in Fig. 8(b), which shows that it resembles the original Schwarzschild case but differs near the singularities. Moreover, for purposes of comparison, for some large $M$ cases, we have drawn the angular components of geodesic deviation in Fig. 9, which shows the angular separations are not much changed.
It seems appropriate to comment that as it approaches the singularity, the radial and angular components of the geodesic deviation of the GUP-modified Schwarzschild black hole are stretched steeper and larger than the ones of the original Schwarzschild black hole, respectively. It looks peculiar since we usually expect quantum gravity effects such as a GUP to make physical quantities more finite, not worse. However, there is also an example of applying GUP to white dwarfs, which leads to a peculiar property. A GUP can remove the Chandrasekhar limit Rashidi:2015rro ; Ong:2018zqn ; Mathew:2020wnx , and thus white dwarfs can be seemingly arbitrarily large. In this respect, a GUP correction can lead to more divergent behaviors for some physical quantities, which goes against our usual expectations and should be further studied.
@@ VI Discussion
Finally, it seems appropriate to comment that our coordinate systems still contains a singularity at $r=\sqrt{\alpha}$, implying that the geodesics is still incomplete inside the black hole. In this respect, it would be interesting to construct and analyze the tidal effect, related to the works Dvali:2010bf ; Isi:2013cxa ; Li:2016yfd ; Alencar:2023wyf , on GUP-induced geodesically compete non-singular black holes through further investigation.